OpenStudy (anonymous):

Simplify 3√123 Answers: 192√2 8√2 11√2 24√2

OpenStudy (angela210793):

none seems right to me...

OpenStudy (saifoo.khan):

Righto

OpenStudy (anonymous):

agree.... i m a kindergarten student

OpenStudy (anonymous):

can someone help me please.

OpenStudy (saifoo.khan):

lol, your options are wrong.

OpenStudy (saifoo.khan):

\[\sqrt{123}\]

OpenStudy (saifoo.khan):

is its like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

saifoo online a

OpenStudy (lalaly):

none of them is right

OpenStudy (saifoo.khan):

\[\Huge \sqrt[3]{123}\] \[\Huge 3\sqrt{123}\] which one?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

is 123 a prime number

OpenStudy (anonymous):

safi kahan hai??????????????

OpenStudy (anonymous):

i am so sorry its 128

OpenStudy (anonymous):

ok

OpenStudy (angela210793):

=24sqr2

OpenStudy (saifoo.khan):

\[\Huge \color\red {24\sqrt{2}}\]

OpenStudy (anonymous):

4 th is correct

OpenStudy (angela210793):

3sqr128=3sqr(64*2)=8*3sqr2=24sqr2

OpenStudy (anonymous):

angela to the rescue